Operations Report – January 26th

Crew 203 Operations Report 26-JAN-2019

SOL: 14 (The final one for us here)

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): 15hr 55min; Turned on last night (25Jan2019) at 16:45; Turned off this morning (26Jan2019) at 8:40; Turned on today (26Jan2019) at 18:55.

Solar SOC – Turned on (25Jan2019) 91%; Turned off (26Jan2019) 100%; Turned on (26Jan2019) 85%

Diesel Reading – Under 1/2 tank.

Propane Reading – Over 79%

Ethanol Free Gasoline – Nothing to report.

Water (auxiliary tank) – Not in use.

Water (static tank) – About 25%; 130 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 29%; 85 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 8:00, 17:05 and 19:10 to refill tank

Water Meter: 01405365

Toilet tank emptied: No.

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used.

Hours: 70.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Not used.

Hours: 46.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used.

Hours: 72.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Used to get crew to town for Hamburgers 🙂

CrewCar used and why, where? Used to get crew 204 from Grand Junction to MDRS

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Suit #10 and Suit #6 batteries still don´t charge.

Summary of Hab operations: It was a very hard cleaning day with a delicious Hamburger as reward. Nothing new to report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report.

Our last day on Mars, it was a real pleasure for me to be here as part of crew 203, thanks for the support, I hope my work has contributed to the program. Thanks to all.

Freddy Castaneda

PS. A special greeting to my daughter, I’ll see you soon. 🙂